# Upper Approximation Bounds for Neural Oscillators

### v1

@@ 3.2 Second-Order Dynamical Systems
Remark 11: Theorem 2 demonstrates that when the target causal continuous operator can be explicitly represented as a uniformly asymptotically incrementally stable second-order dynamical system, the widths of the hidden layers in the two MLPs utilized in the neural oscillator scale polynomially with the reciprocals of the approximation errors. This implies that the curse of parametric complexity of the neural oscillator in this situation is mitigated. An analogous approximation bound for RNNs modeling the first-order dynamical systems has been obtained in Hanson and Raginsky (2020).
@@
Remark 12: Approximating an ODE system by a neural ODE involves approximating the derivative function of the target ODE using an MLP. The approximation bound for an MLP approximating the derivative function is established over a compact domain of the derivative function’s arguments and the MLP’s arguments. The derivative function’s arguments are input functions and the solutions of the target ODE. The MLP’s arguments are input functions and the solutions of the neural ODE. Therefore, in establishing an approximation bound for a neural ODE approximating an ODE system, where the derivative function of the target ODE is approximated by the MLP of the neural ODE, one critical issue is ensuring that the values of all possible solutions from the neural ODE, driven by a set of considered input functions, can lie within the compact domain, where the error between the MLP of the neural ODE and the target derivative function can be quantified. To meet this requirement, Lemma 7 in Hanson and Raginsky (2020) proposes to approximate the derivative function of a target ODE using a compactly supported function, such that the values of the solutions from the resulting neural ODE, whose derivative function is modeled by this compactly supported function, can be bounded in a compact domain. In this study, the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs are established in Eqs. (27) and (28) of Lemma 6. Based on the result of Eq. (28), the error between the MLP ${\it{\Gamma}}(\cdot)$ and the target derivative function $g(\cdot)$ in Eq. (29) is quantified over the compact domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$. And all possible values of the solutions $\mathbf{x}(t)$ and ${\mathbf{x}}^{\prime}(t)$ to the second-order ODE in Eq. (1) with ${\it{\Gamma}}(\cdot)$ driven by all $\mathbf{u}\in K$ are within the domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$, as highlighted in Eq. (99), Eq. (100), and Eq. (106) in the proof of Theorem 2 in B.3.
4 Numerical Study
@@
In this section, a Bouc-Wen nonlinear system driven by stochastic excitations is employed to validate the results in Theorem 1 and Theorem 2. A numerical study in Huang and Xia (2025) has shown that the neural oscillator can effectively learn the mapping relationship between the input excitation and the output response of a Bouc-Wen system exhibiting plastic deformation. The differential equation of a 5-degree of freedom Bouc-Wen system is
$\left\{\begin{aligned} &\mathbf{M}\ddot{\mathbf{X}}(t)+\mathbf{C}\dot{\mathbf{X}}(t)+\lambda{\mathbf{K}\mathbf{X}}(t)+(1-\lambda){\tilde{\mathbf{K}}\mathbf{Z}}(t)=-{\mathbf{M}}{U_{\text{e}}}(t)\\ &{{\dot{Z}}_{i}}(t)={\dot{\tilde{X}}}_{i}(t)-\beta\left|{\dot{\tilde{X}}_{i}}(t)\right|{\left|{Z_{i}}(t)\right|^{s-1}}{Z_{i}}(t)-\gamma{\dot{\tilde{X}}_{i}}(t){\left|{{Z_{i}}(t)}\right|^{s}},\;\;i=1,2,\ldots,5\end{aligned}\right$., (33)
where ${\mathbf{X}}=\left[X_{1}(t),X_{2}(t),\dots,X_{5}(t)\right]^{\top}$, $\tilde{\bf{X}}(t)=[X_{1}(t),X_{2}(t)-X_{1}(t),$ $X_{3}(t)-X_{2}(t),\dots,X_{5}(t)-X_{4}(t)]^{\top}$, ${\mathbf{Z}}=\left[Z_{1}(t),Z_{2}(t),\dots,Z_{5}(t)\right]^{\top}$, ${\mathbf{X}}(0)={\mathbf{X}}^{\prime}(0)={\mathbf{Z}}(0)={\mathbf{0}}$,
@@
$m=1382.4$ kilogram, $k=1.7\times 10^{6}$ Newton/meter, $\mathbf{C}$ is determined using $\mathbf{M}$ and $\mathbf{K}$ such that the damping ratios of all response modes are 0.05, $\lambda=0.01$, $\beta=2$, $\gamma=2$, and $s=3$. $U_{\text{e}}(t)$ in Eq. (33) represents different stochastic excitations in the following two cases. The ODE system in Eq. (33) is solved numerically using a fourth-order Runge-Kutta method (Griffiths and Higham, 2010) with a time discretization of $\Delta t=0.01$ seconds over the time interval $[0,10]$. For the numerical implementation of training, the neural oscillator is discretized using a second-order Runge–Kutta scheme (Huang and Xia, 2025). The approach for training the neural oscillator with input-output pair samples is briefly described in E. All code for this numerical study is available at: https://github.com/ZifengH22/Upper-Approximation-Bounds-for-Neural-Oscillators.
4.1 Case 1
@@
where $c_{j},j=1,2,\ldots,5$ are are independent random variables uniformly distributed on the interval $\left[-35,35\right]$. $5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$, where $i=0,1,2,\dots,999$ and $l=1,2,\dots,5\times 10^{4}$, are simulated. Their induced $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of the extreme value process $E_{X_{5}}(t)=\max_{\tau\in\left[0,t\right]}\left|X_{5}(\tau)\right|$ are computed by numerically solving Eq. (33). The mapping from $U_{\text{e}}(t)$ to $E_{X_{5}}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$. The decay rate in Theorem 1 is validated using this case.
@@
In this case, five neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\ldots,5$, are considered. The sizes of the five one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ are the same. Their widths of input, hidden, and output layers are 11, 10, and 5, respectively. The activation functions of the five MLPs ${\it{\Gamma}}_{i}(\cdot)$ are $\sigma_{\mathrm{ReLU}}(\cdot)$. The five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ corresponding to the five neural oscillators share the same input, hidden, and output layer widths of 7, 10, and 1, respectively. The numbers of hidden layers in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$ are 1, 2, 3, 4, and 5, respectively. The activation functions of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ are the Parametric Rectified Linear Unit (PReLU) $\sigma_{\text{PReLU}}(x)=\max(0,x)+\alpha\min(0,x)$, where $\alpha$ is a parameter learned during training (He et al., 2015). Since $\sigma_{\mathrm{ReLU}}(\cdot)$ is a special case of $\sigma_{\text{PReLU}}(\cdot)$, the result in Theorem 1 also holds for the neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$ when ${{\it{\Pi}}}_{i}(\cdot)$ employs $\sigma_{\text{PReLU}}(\cdot)$.
@@
The $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of output $E_{X_{5}}(t)$ are arranged in a descending order with respect to their maximum values $\hat{E}_{X_{5},l}=\max_{t_{i}\in\left[0,10\right]}E_{X_{5},l}(t_{i})$. Then, 10000 samples $E_{X_{5},l}(t_{i})$, whose maximum values $\hat{E}_{X_{5},l}$ are approximately equally spaced from the largest one to the smallest one of the $5\times 10^{4}$ $\hat{E}_{X_{5},l}$, along with their corresponding input samples $U_{\text{e},l}(t_{i})$, are selected. Among these 10000 sample pairs, the $5^{\text{th}}$, $10^{\text{th}}$, $15^{\text{th}}$$\dots,$ $10000^{\text{th}}$ pairs (total 2000 pairs) are selected as a validation dataset, and the remaining 8000 sample pairs constitute a training dataset. As this study is concerned solely with validating the approximation error bound rather than investigating the generalization performance, this sampling strategy is appropriate to build the training data. Each of the five neural oscillators is trained using the training approach outlined in E with the 8000 sample pairs. The 8-th power loss function $\ell_{8}$ in Eq. (118) is employed to ensure that the neural oscillator adequately learns the target mapping with respect to the $L^{\infty}$-norm. The training phase comprises 10000 epochs, with a batch size of 2000 for each epoch. The initial learn rate is 0.005. The learning rate drop period is 100 epochs. The learning rate drop factor is 0.9772. To mitigate the risk of exploding gradients during training, the gradient norm clipping strategy (Pascanu et al., 2013) with a threshold of 1 is employed for the five MLPs ${\it{\Gamma}}_{i}(\cdot)$. For the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, their norm clipping thresholds are $1,\sqrt{203/92},\sqrt{314/92},\sqrt{425/92},\sqrt{536/92}$, respectively, where $92,203,314,425,$ and $536$ are the numbers of learnable parameters in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$. Batch normalization (Ioffe and Szegedy, 2015) is applied to the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ to facilitate training due to the depth of their hidden layers. Since batch normalization only adds an affine transformation to the input of each hidden layer of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, the result of Theorem 1 still holds for ${{\it{\Pi}}}_{i}(\cdot)$ with batch normalization. For the first neural oscillator with ${\it{\Gamma}}_{1}(\cdot)$ and ${{\it{\Pi}}}_{1}(\cdot)$, the initial values of the learnable network parameters are randomly selected. For the second neural oscillator with ${\it{\Gamma}}_{2}(\cdot)$ and ${{\it{\Pi}}}_{2}(\cdot)$, the initial parameter values of ${\it{\Gamma}}_{2}(\cdot)$ are taken directly from the parameters of ${\it{\Gamma}}_{1}(\cdot)$ after the last update. The initial parameter values for the first hidden layer and output layer of ${{\it{\Pi}}}_{2}(\cdot)$ are taken directly from the parameters of the corresponding layers in ${{\it{\Pi}}}_{1}(\cdot)$ after the last update. The second hidden layer of ${{\it{\Pi}}}_{2}(\cdot)$ is initialized as an identity mapping. The MLPs in the remaining three neural oscillators are initialized in the same manner. For each neural oscillator, the neural network achieving the smallest validation loss under the $L^{\infty}$-norm is selected as the final model.
@@
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{E}_{X_{5},l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors under $L^{\infty}$-norm and $L^{2}$-norm, which are respectively denoted as $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$, are computed as
@@
$\tilde{\varepsilon}_{E_{X},\infty}=\frac{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|\tilde{E}_{X_{5},l}(t_{i})-E_{X_{5},l}(t_{i})\right|}{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|E_{E_{5},l}(t_{i})\right|}$ (36)

### v2

@@ 3.2 Second-Order Dynamical Systems
Remark 12: Theorem 2 demonstrates that when the target causal continuous operator can be explicitly represented as a uniformly asymptotically incrementally stable second-order dynamical system, the widths of the hidden layers in the two MLPs utilized in the neural oscillator scale polynomially with the reciprocals of the approximation errors under the $\it{L^{\infty}}$-norm. This implies that the curse of parametric complexity of the neural oscillator in this situation is overcome, in contrast to the result established in Theorem 1 that is discussed in Remark 9. The distinction stems from the fact the target operator in Theorem 2 is an explicit second-order ODE and the neural oscillator only needs to use its MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ to approximate the functions $g(\cdot)$ and $h(\cdot)$, respectively. Since the argument dimensions of $g(\cdot)$ and $h(\cdot)$ are fixed hyperparameters for a given target ODE, the neural oscillator can achieve polynomial convergence rates independent of the dimension of the function set $K$ containing $\mathbf{u}(t)$.
Remark 13: The dimension-independent convergence rates of 0.5-0.5 in Eqs. (31) and (32) are achieved because $g(\cdot)$ and $h(\cdot)$ in Theorem 2 admit Barron function extensions over any compact domain of their respective arguments, with the corresponding spectral Barron norms defined in Theorem 4 of Siegel and Xu (2023) being uniformly bounded. Theorem 2 in Barron (1992) indicates that similar error bounds, with the convergence rates of 0.5-0.5 as in Eqs. (30)–(32), can also be achieved when the tanh or sigmoid activation function is utilized. $g(\cdot)$ and $h(\cdot)$ belong to the variation space generated by a spectral dictionary that was established by Siegel and Xu (2023). The Barron spectrum space ${\mathcal{B}_{1,2}}(\mathbb{R}^{r})$ in Definition 2.1 of Meng and Ming (2022) is a subset of the possible $h(\cdot)$ in the variation space. Theorem 4.3 in Meng and Ming (2022) demonstrates that the Sobolev space ${\mathcal{W}^{\alpha,2}}(\mathbb{R}^{r})$ satisfying $\alpha>1+r/2>0$ is embedded in ${\mathcal{B}_{1,2}}(\mathbb{R}^{r})$. Therefore, ${\mathcal{W}^{\alpha,2}}(\mathbb{R}^{r})$ with $\alpha>1+r/2>0$ is also a subset of the possible $h(\cdot)$. The approximation error in Theorem 2 has two parts $O\left(w_{{\it{\Gamma}}}^{-1/2}\right)+O\left(w_{{\it{\Pi}}}^{-1/2}\right)$. For the second one, it arises from approximating $h(\cdot)$ using a one-hidden-layer ReLU MLP ${\it{\Pi}}(\cdot)$. Since ${\mathcal{W}^{\alpha,2}}(\mathbb{R}^{r})$ is a subset of the possible $h(\cdot)$, and the one-hidden-layer ReLU MLP ${\it{\Pi}}(\cdot)$ for approximating ${\mathcal{W}^{\alpha,2}}(\mathbb{R}^{r})$ has a lower bound in the form of ${\it{\Omega}}\left[w_{\it{\Pi}}^{-\alpha/(r-1)}\right]$ (Maiorov, 1999). Thus, $\left(2+0.5r\right)/\left(r-1\right)$ is a lower-bound convergence rate which approaches the dimension-independent rate of 0.5-0.5 in Theorem 2 as the argument dimension $r$ of $h(\cdot)$ increases. For the first part of the approximation error in Theorem 2, since two ODEs with two different derivative term functions, e.g., $g_{1}\left[{{\hat{\bf x}}(t),{\hat{\bf x}}^{\prime}(t),{\bf{u}}(t)}\right]$ and $g_{2}\left[{{\hat{\bf x}}(t),{\hat{\bf x}}^{\prime}(t),{\bf{u}}(t)}\right]$, can have the same solution $\hat{\mathbf{x}}(t)$ under the same input $\mathbf{u}(t)$, as long as $g_{1}(\cdot)$ and $g_{2}(\cdot)$ agree on the trajectories of ${\hat{\bf x}}(t),{\hat{\bf x}}^{\prime}(t)$, and ${\bf{u}}(t)$, a lower bound of the MLP ${\it{\Gamma}}(\cdot)$ to approximate the function $g(\cdot)$ in Eq. (29) is not sufficient to establish a lower bound of the neural ODE in the neural oscillator to approximate the target ODE. Currently, research on the lower bounds for the neural ODEs viewed as mappings from input functions to solution functions remains limited. Bittner (2023) established the Vapnik-Chervonenkis (VC) dimension bounds for discrete-time RNNs and liquid time constant networks. Since a neural oscillator with a fixed-step ODE solver is structurally similar to these two discrete-time neural networks, extending their framework to derive the VC dimension of the neural oscillator and subsequently establishing its lower bound represents a natural direction for future research.
@@
Remark 14: Approximating an ODE system by a neural ODE involves approximating the derivative function of the target ODE using an MLP. The approximation bound for an MLP approximating the derivative function is established over a compact domain of the derivative function’s arguments and the MLP’s arguments. The derivative function’s arguments are the input functions and solutions of the target ODE. The MLP’s arguments are the input functions and solutions of the neural ODE. Therefore, in establishing an approximation bound for a neural ODE approximating an ODE system, where the derivative function of the target ODE is approximated by the MLP of the neural ODE, one critical issue is ensuring that the values of all possible solutions from the neural ODE, driven by a set of considered input functions, can lie within the compact domain, where the error between the MLP of the neural ODE and the target derivative function can be quantified. To meet this requirement, Lemma 7 in Hanson and Raginsky (2020) proposes to approximate the derivative function of a target ODE using a compactly supported function, such that the values of the solutions from the resulting neural ODE, whose derivative function is modeled by this compactly supported function, can be bounded in a compact domain. In this study, the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs are established in Eqs. (27) and (28) of Lemma 6. Based on the result of Eq. (28), the error between the MLP ${\it{\Gamma}}(\cdot)$ and the target derivative function $g(\cdot)$ in Eq. (29) is quantified over the compact domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$. And all possible values of the solutions $\mathbf{x}(t)$ and ${\mathbf{x}}^{\prime}(t)$ to the second-order ODE in Eq. (1) with ${\it{\Gamma}}(\cdot)$ driven by all $\mathbf{u}\in K$ are within the domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$, as highlighted in Eqs. (104), (105), and (110) in the proof of Theorem 2 in B.3.
Remark 15: The approximation error bounds established in Theorems 1 and 2 are existence results. These results show that there exist weight configurations for the two MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ such that the neural oscillator in Eq. (1) can approximate the target causal operators or dynamical systems within the established error bounds. Furthermore, upper bounds on the generalization errors of the neural oscillator, accounting for the error induced by the finite number of training samples, have been established in Huang et al. (2026). However, these theoretical results do not address the training problem. Specifically, they do not specify which training algorithms can efficiently identify models that can achieve the theoretically predicted approximation and generalization error bounds. This gap between theory and practice in neural networks is a frontier research challenge. Notably, Kratsios et al. (2025) proposed a specific neural network architecture that combines the attention mechanism with a random neural network, for which the theoretically predicted generalization performance can be achieved by trained models with high probability using a finite number of samples. Developing specialized neural oscillator architectures, together with efficient training strategies, that enable trained models to realize the theoretically predicted error bounds, is a promising direction for future research.
Remark 16: Since the error bounds established in Theorems 1 and 2 contain several constants that are difficult to evaluate in practice, they are not directly suitable for determining appropriate neural network sizes for specific problems under prescribed error tolerances. Nevertheless, these theoretical results still provide useful guidance. In practice, initial network sizes can be selected based on prior experience. Then, by combining the convergence rates derived in Theorems 1 and 2 with the approximation errors obtained from the chosen network configurations, the network sizes can be further estimated and adjusted to achieve the desired accuracy. Furthermore, the proof process of the two theorems provides the mathematical explanations for the possible behaviors of the neural oscillator in approximating various operator mappings. This theoretical understanding of the neural oscillator can promote its more effective application compared with other models. For example, the time-continuous RNN, consisting of a neural ODE with a linear readout function, is a universal approximator for dynamical ODE mappings (Hanson et al., 2021). However, for causal mappings between temporal function spaces that cannot be expressed as dynamical ODEs, the approximation capability of time-continuous RNNs remains unclear. In contrast, Theorem 1 in this study demonstrates that the neural oscillator can universally approximate causal and continuous operators, including those that cannot be represented in the form of dynamical ODEs. While attention-based transformers can serve as universal approximators for continuous operators between continuous function spaces (Kovachki et al., 2023), they may require large network sizes when applied to target mappings that admit explicit ODE forms due to architectural mismatch. In contrast, Theorem 2 in this study demonstrates that the neural oscillator can directly approximate the derivative terms of the target ODEs, achieving the desired approximation results with more effective network sizes.
4 Numerical Studies
In this section, four numerical cases are presented to validate the convergence rates of the derived approximation error bounds and to illustrate the convergence properties of the neural oscillator. The code for all the numerical cases is available at: https://github.com/ZifengH22/Upper-Approximation-Bounds-for-Neural-Oscillators.
4.1 Case 1
@@
In Cases 1 and 2, a Bouc-Wen nonlinear system driven by stochastic excitations is considered. A numerical study in Huang and Xia (2025) has shown that the neural oscillator can effectively learn the mapping relationship between the input excitation and the output response of a Bouc-Wen system exhibiting plastic deformation. The differential equation of a 5-degree-of-freedom Bouc-Wen system is
[…]
where ${\mathbf{X}}=\left[X_{1}(t),X_{2}(t),\dots,X_{5}(t)\right]^{\top}$, $\tilde{\bf{X}}(t)=[X_{1}(t),X_{2}(t)-X_{1}(t),\dots,X_{5}(t)-X_{4}(t)]^{\top}$, ${\mathbf{Z}}=\left[Z_{1}(t),Z_{2}(t),\dots,Z_{5}(t)\right]^{\top}$, ${\mathbf{X}}(0)={\mathbf{X}}^{\prime}(0)={\mathbf{Z}}(0)={\mathbf{0}}$,
@@
$m=1382.4$ kilogram, $k=1.7\times 10^{6}$ Newton/meter, $\mathbf{C}$ is determined using $\mathbf{M}$ and $\mathbf{K}$ such that the damping ratios of all response modes are 0.05, $\lambda=0.01$, $\beta=2$, $\gamma=2$, and $s=3$. $U_{\text{e}}(t)$ in Eq. (33) represents different stochastic excitations in Cases 1 and 2. The ODE system in Eq. (33) is solved numerically using a fourth-order Runge-Kutta method (Griffiths and Higham, 2010) with a time discretization of $\Delta t=0.01$ seconds over the time interval $[0,10]$. For the numerical implementation of training, the neural oscillator is discretized using a second-order Runge–Kutta scheme (Huang and Xia, 2025). The approach for training the neural oscillator with input-output pair samples is briefly described in E.
@@
where $c_{j},j=1,2,\ldots,5$ are are independent random variables uniformly distributed on the interval $\left[-35,35\right]$. $5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$, where $i=0,1,2,\dots,999$ and $l=1,2,\dots,5\times 10^{4}$, are simulated. Their induced $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of the extreme value process $E_{X_{5}}(t)=\max_{\tau\in\left[0,t\right]}\left|X_{5}(\tau)\right|$ are computed by numerically solving Eq. (33). The mapping from $U_{\text{e}}(t)$ to $E_{X_{5}}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$.
@@
In this case, five neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\ldots,5$, are utilized to validate the convergence rate in Theorem 1. The sizes of the five one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ are the same. Their input, hidden, and output layer widths are 11, 10, and 5, respectively. The activation functions of the five MLPs ${\it{\Gamma}}_{i}(\cdot)$ are $\sigma_{\mathrm{ReLU}}(\cdot)$. The five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ corresponding to the five neural oscillators share the same input, hidden, and output layer widths of 7, 10, and 1, respectively. The numbers of hidden layers in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$ are 1, 2, 3, 4, and 5, respectively. The activation functions of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ are the PReLU function $\sigma_{\text{PReLU}}(x)=\max(0,x)+\alpha\min(0,x)$, where $\alpha$ is a parameter learned during training (He et al., 2015). Since $\sigma_{\mathrm{ReLU}}(\cdot)$ is a special case of $\sigma_{\text{PReLU}}(\cdot)$, the result in Theorem 1 also holds for the neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$ when ${{\it{\Pi}}}_{i}(\cdot)$ employs $\sigma_{\text{PReLU}}(\cdot)$.
@@
The $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of output $E_{X_{5}}(t)$ are arranged in a descending order with respect to their maximum values $\hat{E}_{X_{5},l}=\max_{t_{i}\in\left[0,10\right]}E_{X_{5},l}(t_{i})$. Then, 10000 samples $E_{X_{5},l}(t_{i})$, whose maximum values $\hat{E}_{X_{5},l}$ are approximately equally spaced from the largest one to the smallest one of the $5\times 10^{4}$ $\hat{E}_{X_{5},l}$, along with their corresponding input samples $U_{\text{e},l}(t_{i})$, are selected. Among these 10000 sample pairs, the $5^{\text{th}}$, $10^{\text{th}}$, $15^{\text{th}}$$\dots,$ $10000^{\text{th}}$ pairs (total 2000 pairs) are selected as a validation dataset, and the remaining 8000 sample pairs constitute a training dataset. As this study is concerned solely with validating the approximation error bound rather than investigating the generalization performance, this sampling strategy is appropriate to build the training data. Each of the five neural oscillators is trained using the training approach outlined in E with the 8000 sample pairs. The $8^{\text{th}}$ power loss function $\ell_{8}$ in Eq. (122) is employed to ensure that the neural oscillator adequately learns the target mapping with respect to the $L^{\infty}$-norm. The training phase comprises 10000 epochs, with a batch size of 2000 for each epoch. The initial learn rate is 0.005. The learning rate drop period is 100 epochs. The learning rate drop factor is 0.9772. To mitigate the risk of exploding gradients during training, the gradient norm clipping strategy (Pascanu et al., 2013) with a threshold of 1 is employed for the five MLPs ${\it{\Gamma}}_{i}(\cdot)$. For the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, their norm clipping thresholds are $1,\sqrt{203/92},\sqrt{314/92},\sqrt{425/92},\sqrt{536/92}$, respectively, where $92,203,314,425,$ and $536$ are the numbers of learnable parameters in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$. Batch normalization (Ioffe and Szegedy, 2015) is applied to the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ to facilitate training due to the depth of their hidden layers. Since batch normalization only adds an affine transformation to the input of each hidden layer of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, the result of Theorem 1 still holds for ${{\it{\Pi}}}_{i}(\cdot)$ with batch normalization. For the first neural oscillator with ${\it{\Gamma}}_{1}(\cdot)$ and ${{\it{\Pi}}}_{1}(\cdot)$, the initial values of the learnable network parameters are randomly selected. For the second neural oscillator with ${\it{\Gamma}}_{2}(\cdot)$ and ${{\it{\Pi}}}_{2}(\cdot)$, the initial parameter values of ${\it{\Gamma}}_{2}(\cdot)$ are taken directly from the parameters of ${\it{\Gamma}}_{1}(\cdot)$ after the last update. The initial parameter values for the first hidden layer and output layer of ${{\it{\Pi}}}_{2}(\cdot)$ are taken directly from the parameters of the corresponding layers in ${{\it{\Pi}}}_{1}(\cdot)$ after the last update. The second hidden layer of ${{\it{\Pi}}}_{2}(\cdot)$ is initialized as an identity mapping. The MLPs in the remaining three neural oscillators are initialized in the same manner. For each neural oscillator, the neural network achieving the smallest validation loss under the $L^{\infty}$-norm is selected as the final model.
Refer to caption(a)
Refer to caption(b)
Figure 2: Approximation errors $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ versus the number of hidden layers $H_{{\it{\Pi}}_{i}}$ of ${\it{\Pi}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{E_{X},\infty}$ versus $H_{{\it{\Pi}}_{i}}$. (b) $\tilde{\varepsilon}_{E_{X},2}$ versus $H_{{\it{\Pi}}_{i}}$.
@@
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{E}_{X_{5},l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors under the $L^{\infty}$-norm and $L^{2}$-norm, which are respectively denoted as $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$, are computed as
@@
$\tilde{\varepsilon}_{E_{X},\infty}=\frac{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|\tilde{E}_{X_{5},l}(t_{i})-E_{X_{5},l}(t_{i})\right|}{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{X_{5},l}}(t_{i})\right|}$ (36)